A study is conducted to investigate the impact of depression on a patient’s ability to survive cardiac disease. They evaluated 477 participants for depression and followed for 4 years. There were 374 patients with no depression (ND) and 69 of them died. Out of 103 patients with minor or major depression (D), 28 died. Does it appear that the depressed (D) patients with cardiac disease are more likely to die than non-depressed (ND) patients with cardiac disease? Conduct a test of hypothesis at 5% level of significance. What is the p-value for the test?
Answer #1

x1 = 69

n1 = 374

x2 = 28

n2 = 103

pl = 1 = 2 = 0.1845 p1 - 69 .

p2 = 100 = 0.2718

Claim: The depressed (D) patients with cardiac disease are more likely to die than non-depressed (ND) patients with cardiac disease

The null and alternative hypothesis is

HO: P1 < P2

Hl: Pl>P2

Level of significance = 0.05

Test statistic is

pl - p2 Vp* (1-P) (1 + m2)

01 + x2 P nl + n2

69 + 28 P = 374 + 103 = 0.2034

10.1845 -0.2718 (0.2034 *(1-0.20341) +( + )

-0.0874 0.1620 * 0.0124 = =-1.95

P-value = P(Z < -1.95) = 0.0256

P-value < 0.05 we reject null hypothesis.

Conclusion:

The depressed (D) patients with cardiac disease are more likely to die than non-depressed (ND) patients with cardiac disease.